OpenStudy (mrdoodbromanguy):

A jar contains 12 nickels, 12 dimes, and 12 quarters. A coin is chosen at random from the jar. What is the probability that the coin chosen is a quarter? A. 1/36 B. 1/12 C. 1/4 D. 1/3

OpenStudy (aveline):

There are a total of 36 coins. We have 12 quarters, so the chance is 12/36, which simplifies to 1/3

OpenStudy (bmk614):

Quarters/Total=12/36=1/3

OpenStudy (mrdoodbromanguy):

idk who to medal :D